# High-dimensional dynamics in low-dimensional networks

### v2

@@ Discussion
Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics. In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Supplementary Text S.1 for a more complete comparison). Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
Low-rank suppression might seem obvious in hindsight: When $W_{0}$ is normal and symmetric, the combination of stability and large singular values implies eigenvalues with large, negative real part, so dynamics are highly suppressive in one direction. Similarly, high-dimensional dynamics might seem like an obvious consequence of high-dimensional inputs. However, the example in Figure 2 coupled with the results in Claims 1 and 2 show that reality is more nuanced: Large negative eigenvalues are not necessary for low-rank suppression, and high-dimensional input is not sufficient for high-dimensional responses.
@@ Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations – except for Figure 6 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
@@ Acknowledgments
Supplementary Materials
Figure S.1: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure S.2: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure S.3: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
Figure S.4: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor (c=2c=-2), and the dynamics obey Eq. (9). The instability produced by taking $\rho>1$ generates chaotic dynamics [27, 3].
Figure S.5: Dynamics of an unstable, non-chaotic network with rank-one structure. Same as Figure 1 except $c=10$ and the dynamics obey Eq. (9). The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
Figure S.6: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (9), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure S.7: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
Figure S.8: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
Figure S.9: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
Figure S.10: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.
S.1 Comparison to previous work on dynamics of low-rank networks
@@
Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
@@
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2 and Supplementary Figure S.3 which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Supplementary Figure S.11 and S.12).
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Supplementary Figure S.11). Similar examples and results are considered in [10]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Supplementary Figure S.12), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Supplementary Figure S.12d; compare to Figure 2 and Supplementary Figure S.3).
@@
S.2 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
@@
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Supplementary Figure S.2), and would not be related to the intrinsic dynamics of interactions within the network.
@@
Figure S.13: Comparison between theory and simulations for shorter and longer simulations. a) Same as Figure 1d except we added the theoretical values form Eq. (16) (red dots) and the values obtained from Eq. (17) (green dots). The network was simulated for $T=5\times 10^{3}$ time units. b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (16) as red dots (Supplementary Figure S.13a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (15) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (16) would be replaced by
@@
is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (17) gives a much more accurate prediction (Supplementary Figure S.13a, green dots). This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
@@
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (15). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (13) instead of the identity itself, as in Eq. (15). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$). Next, in Supplementary Text S.3, we generalize away from the assumption of slow inputs using a Laplace transform.
S.3 The analysis of fast and transient dynamics.
@@
Note also that Eq. (13) from Supplementary Text S.2 can be derived from Eqs. (20) and (19) above by first taking $s=-2\pi if$ in Eq. (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the approach in this section is a generalization of the approach in Supplementary Text S.2.

### v4

@@ Discussion
Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics under some assumptions. In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Appendix A for a more complete comparison). Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
Our results also rely on a linearization of the dynamics around a stable fixed point. The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question. However, nonlinear networks can produce any dynamics [29, 30] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks. Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then using bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations. Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d). A large, negative eigenvalue is well-known to produce a suppressive dynamic. However, the results in Figure 2, Figure 12 in Appendix D, and Claim 1 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression. Low-rank suppression occurs in strongly low-rank networks regardless of their eigenvalue spectrum.
@@ Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations – except for Figure 6 and Figure 19 in Appendix D – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Figure 19 in Appendix D were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
@@ Acknowledgments
Appendix
Appendix A Comparison to previous work on dynamics of low-rank networks
@@
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 12 in Appendix D which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Figures 7 and 8 in Appendix D).
Figure 7: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure 8: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 7). Similar examples and results are considered in [10]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 8), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 8d; compare to Figure 2 and Figure 12).
@@
Appendix B Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
@@
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Figure 11), and would not be related to the intrinsic dynamics of interactions within the network.
@@
Figure 9: Comparison between theory and simulations for shorter and longer simulations. a) Same as Figure 1d except we added the theoretical values form Eq. (16) (red dots) and the values obtained from Eq. (17) (green dots). The network was simulated for $T=5\times 10^{3}$ time units. b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (16) as red dots (Figure 9a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (15) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (16) would be replaced by
@@
is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (17) gives a much more accurate prediction (Figure 9a, green dots). This confirms that the error in the red dots from Figure 9a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (16) was more accurate (Figure 9b, red dots), further confirming that the errors in Figure 9a are due largely to sampling error.
@@
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (15). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (13) instead of the identity itself, as in Eq. (15). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$). Next, in Appendix C, we generalize away from the assumption of slow inputs using a Laplace transform.
Appendix C The analysis of fast and transient dynamics.
@@
Note also that Eq. (13) from Appendix B can be derived from Eqs. (20) and (19) above by first taking $s=-2\pi if$ in Eq. (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the approach in this section is a generalization of the approach in Appendix B.
@@
Appendix D Additional Simulations and Data
This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text.
Figure 10: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure 11: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure 12: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
Figure 13: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$ and the magnitude of $c$ was increased by the same factor (c=40c=-40) to maintain the same degree of low-rank structure (the same ratio between the spectral radius of the random part, $\rho$, and the magnitude of the low-rank part, $|c|$), and the dynamics obey Eq. (9). The instability produced by taking $\rho>1$ generates chaotic dynamics.
Figure 14: Dynamics of an unstable, non-chaotic network with rank-one structure. Same as Figure 1 except $c=10$ and the dynamics obey Eq. (9). The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
Figure 15: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (9), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure 16: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
Figure 17: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
Figure 18: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.